OpenStudy (anonymous):

Which of the following represents a perfect square trinomial? a x2 – 14x + 49 b x2 + 14x + 49 c 4x2 + 28x + 49 d All of the above

OpenStudy (anonymous):

b (x+7)(x+7)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

sorry all the above a= (x-7)^2 b=(x+7)^2 c= (2c+7)^2

OpenStudy (anonymous):

c= (2x+7)^2